An open-top box is constructed from a square sheet of tin, 14 inches on each side, by cutting out equal squares at each corner and then folding up the flaps. Express the volume of the box as a function of x, the length of the side of each cutout square.
Answer by mananth(16946)

x inches length is cut off from the corners
so the base of the box will be (14-x) , (14-x)
and height = x
V= (14-x)^2*x
V=(196-28x+x^2)*x
V=x^3-28x^2+196x